OpenStudy (anonymous):

A rancher has 1200 ft of fencing to construct 4 corrals. Find dimensions of the plot that will maximize the area. PLEASE HELP FINAL EXAM TOMORROW WILL MEDAL AND FAN

OpenStudy (anonymous):

Please help!

OpenStudy (immanuelv):

Area = width*length A(x) = x(1200-2x) A(x) = 1200x - 2x^2 You have a quadratic with a= -2 and b = 1200 Maximum Area occurs where x = -b/2a = -1200/(2*-2) = 300 ft. (width) length =1200-2x = 1200-2*300 = 600 ft. (length) hope that helps

OpenStudy (anonymous):

how did you get x(1200-2x)

OpenStudy (immanuelv):

Let each side perpendicular to the plot be "x". Then the side parallel to the plot is "1200-2x". so you multiply length * width for area

OpenStudy (anonymous):

but how did you get the -2x sorry im confused

OpenStudy (mathstudent55):